Last weekend I drove 330 miles from Los Angeles to Santa Cruz. My car averaged 20.8 miles per
gallon of gasoline during this trip and when I filled up the cost of gasoline was $2.59 (9/10)
a) Use mental math to approximate the total cost of the trip and explain how you arrived at your answer.
b) Write the exact amount of gasoline my car consumed as a fraction.
c) What was the actual cost of the trip? Round your answer to the nearest cent.
1
Expert's answer
2019-12-09T11:35:37-0500

a) The total cost of the trip is:

330/20.8×$2.599 = $41.23.

b) The exact amount of gasoline my car consumed as a fraction is:

330/20.8 = 3300/208 = 825/52.

c) The actual cost of the trip is:

330/20.8×$2.599 = $41.23.